P8-1 (Algo) Explaining the Nature of a Long-Lived Asset and Determining and Recording the Financial Statement Effects of Its Purchase LO8-1, 8-2

[The following information applies to the questions displayed below.]



On January 2, Summers Company received a machine that the company had ordered with an invoice price of $102,000. Freight costs of $720 were paid by the vendor per the sales agreement. The company exchanged the following on January 2 to acquire the machine:

a. Issued 1,900 shares of Summers Company common stock, par value $1 (market value, $3.50 per share).

b. Signed a note payable for $68,000 with an 9.7 percent interest rate (principal plus interest are due April 1 of the current year).

c. The balance of the invoice price was on account with the vendor, to be paid in cash by January 12.

On January 3, Summers Company paid $2,500 cash for installation costs to prepare the machine for use.

P8-1 Part 3

On January 12, Summers Company paid the balance due on its accounts payable to the vendor.

3. Indicate the effects of the purchase and subsequent cash payment on the accounting equation. (Enter decreases to account categories as negative amounts.)

Date Assets = Liabilities + Stockholders’ Equity
Jan 02
Jan 03
Jan 12
Answer #1
Date Assets = Liabilities + Stockholders equity
Jan 02 Equipment 102,000 = Notes Payable 68,000 + Common Stock 1900
Accounts Payable 27,350 Additional paid in capital - Common Stock 4750
Jan 03 Equipment 2500
Cash -2500
Jan 12 Cash -27350 Accounts Payable -27,350

Workings:

Additional paid in capital = 1900 * (3.5 -1) = 4750

Accounts Payable = 102,000 - 68,000 - 1900 - 4750 = 27350